# Diffusion Forcing for Multi-Agent Interaction Sequence Modeling

### v1

@@ 3.2 Latent Motion Encoding
After training the VQ-VAE, we freeze its parameters and use the encoded motion for each person as the token for modeling the joint distribution of multiple people’s motion.
3.3 Multi-Agent Diffusion Forcing Transformer
Having established our per-agent local motion representation via VQ-VAE, we now turn to modeling the joint distribution of multi-agent interactions. Our approach is inspired by the Diffusion Forcing (DF) framework [chen2025diffusion], which applies different noise levels to different tokens in a sequence. We introduce Multi-Agent Diffusion Forcing Transformer (MAGNet), a transformer-based auto-regressive diffusion model trained over sequences of tokens, where each token represents a specific agent at a specific timestep. Each token is informed by each agent’s body pose, shape, and relative temporal transforms and relative transforms to all other agents. Similar to DF, each token receives a different noise level during training. This simple approach enables flexible conditioning over motion history at inference time while preserving temporal coherence between interacting agents. See Figure 3, for illustration of MAGNet, which operates in the latent space, receiving encoded motion tokens and predicting denoised latents that are decoded back to motion trajectories.
Specifically, we define the motion token $\mathbf{m}_{i}^{p}$ for the $p$-th agent ($p\in\{1,\dots,P\}$) at the token timestep $i$ ($i\in\{1,\dots,T/\omega\})$ as the concatenation of three elements:
$\mathbf{m}_{i}^{p}=\left[\mathbf{z}_{i}^{p};\mathbf{T}^{\text{self}\to\text{partner}}_{i\omega\ :(i+1)\omega\ };\Delta\mathbf{T}^{\text{can}}_{i\omega\ :(i+1)\omega\ }\right]\in\mathbb{R}^{D}$ (9)
Where $D$ is the total token dimension. Here $\mathbf{z}^{p}_{i}$, the latent representation of the $p$-th agent’s local body motion equation 5, which is concatenated with the intra-agent temporal transform $\mathbf{\Delta}{T}^{can}_{i\omega\ :(i+1)\omega\ }$ and the pairwise inter-agent transforms $\mathbf{T}^{self\to partner}_{i\omega\ :(i+1)\omega\ }$ across the time window $i\omega,(i+1)\omega)$.
The Diffusion Forcing Transformer (DFoT) processes the complete motion sequence $\mathbf{M}\in\mathbb{R}^{(P\cdot T/\omega)\times D}$ which is constructed by interleaving the tokens of all $P$ agents across all $T^{\prime}=T/\omega$ time steps.
Forward Process. To promote robustness and flexible conditioning, we perturb each clean token $\mathbf{m}_{i}$ with an independent, continuous noise level $\tau_{i}$. We define $\tau_{i}$ as the continuous denoising step, which is independently sampled from a uniform distribution: $\tau_{i}\sim\mathcal{U}(0,1)$. The coefficient $\bar{\alpha}$ is defined using a cosine schedule based on this denoising step $\tau\in[0,1]$$\bar{\alpha}(\tau)=\cos^{2}\left(\frac{\tau+0.008}{1.008}\cdot\frac{\pi}{2}\right).$
The perturbed (noisy) token $\mathbf{m}^{p}_{i}(\tau^{p}_{i})$ is then defined as:
$\mathbf{m}^{p}_{i}(\tau^{p}_{i})=\sqrt{\bar{\alpha}(\tau^{p}_{i})}\mathbf{m}^{p}_{i}+\sqrt{1-\bar{\alpha}(\tau^{p}_{i})}\boldsymbol{\epsilon}^{p}_{i},$ (10)
where $\boldsymbol{\epsilon}^{p}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. For each token $\mathbf{m}^{p}_{i}$, we sample its noise level $\tau^{p}_{i}$ independently across tokens (i.e., i.i.d. over $(i,p)$) and apply the corresponding $\bar{\alpha}(\tau^{p}_{i})$ to perturb $\mathbf{m}^{p}_{i}$.
Transformer Denoiser. The Transformer denoiser $f_{\mathbf{\phi}}$ processes the noised motion token sequence $\mathbf{M}^{(\tau)}$ to predict the clean motion token sequence $\hat{\mathbf{M}}_{\text{0}}$. The noised sequence ($\mathbf{M}^{(\tau)}$) is defined as:
$\mathbf{M}^{(\tau)}=\left[\mathbf{m}_{1}^{1}(\tau_{1}^{1}),\dots,\mathbf{m}_{1}^{P}(\tau_{1}^{P}),\dots,\mathbf{m}^{p}_{T^{\prime}}(\tau^{P}_{T^{\prime}})\right]$ (11)
for $\mathbf{M}^{(\tau)}\in\mathbb{R}^{(P\cdot T^{\prime})\times D}$, where each token $\mathbf{m}^{p}_{i}(\tau^{p}_{i})$ is perturbed by its independent, sampled noise level $\tau^{p}_{i}$.
The denoiser predicts the clean token $\hat{\mathbf{m}}^{p}_{i}$ corresponding to the noisy input $\mathbf{m}^{p}_{i}(\tau^{p}_{i})$. For any given token $(i,p)$ in the sequence, the input embedding $\mathbf{e}^{p}_{i}$ is constructed as: $\mathbf{e}^{p}_{i}=$
$\text{MLP}\left(\left[\mathbf{m}^{p}_{i}(\tau^{p}_{i});\text{SinEmb}(\tau_{i}^{p})\right]\right)+\text{RoPE}(\mathbf{m}^{p}_{i}(\tau_{i}^{p}))+\mathbf{\psi}(p)\text{,}$ (12)
where $\text{SinEmb}(\tau^{p}_{i})$ encodes the noise level, RoPE [su2021roformer] injects the temporal sequencing, and $\mathbf{\psi}(p)$ is a learned positional embedding for agent identity.
The transformer processes the full sequence of these input embeddings $\mathbf{e}_{\text{seq}}=[\mathbf{e}^{1}_{1},\dots,\mathbf{e}^{P}_{T/\omega}]$ to output the predicted clean sequence $\hat{\mathbf{M}}_{\mathbf{0}}$:
$\hat{\mathbf{M}}_{\mathbf{0}}=\mathbf{f}_{\mathbf{\phi}}(\mathbf{M}^{(\tau)},\mathbf{\tau}_{\text{seq}}),$ (13)
where $\mathbf{\tau}_{\text{seq}}=[\tau^{1}_{1},\dots,\tau^{P}_{T/\omega,}]$ is the sequence of sampled noise (matching the order of $\mathbf{M}^{(\tau)}$).
Training Objective. MAGNet is optimized using an $\mathbf{M}_{\mathbf{0}}$-prediction objective via the smooth $\mathbf{L}_{1}$ (Huber) loss across all $N=P\cdot T/\omega$ tokens:
$\mathcal{L}_{\text{Total}}=\mathbb{E}_{\mathbf{M}_{\mathbf{0}},\mathbf{\tau}_{\text{seq}},\boldsymbol{\epsilon}_{\text{seq}}}\left[\|\mathbf{M}_{\mathbf{0}}-\mathbf{f}_{\mathbf{\phi}}(\mathbf{M}^{(\tau)},\mathbf{\tau}_{\text{seq}})\|_{1}\right]$ (14)
The expectation is taken over the ground-truth clean motion sequence $\mathbf{M}_{\mathbf{0}}$, the sequence of sampled noise levels $\mathbf{\tau}_{\text{seq}}$, and the sequence of Gaussian noise samples $\boldsymbol{\epsilon}_{\text{seq}}$.
This loss is decomposed across the individual token components using weighting coefficients $\lambda$:
[…]
where $\mathbf{Z}$ is the ground-truth latent pose, $\mathbf{T}^{\text{self}\to\text{partner}}$ models inter-agent interaction, and $\Delta\mathbf{T}^{\text{can}}$ enforces temporal continuity. All three quantities are defined over the full sequence. The consistency loss, $\mathcal{L}_{\text{c}}$, represents interpersonal velocity consistency and captures the kinematic relationship between agents’ predicted transforms across time:
$\begin{split}\mathcal{L}_{\text{c}}=d_{T}\big(&\mathbf{\hat{T}}^{\text{self}\to\text{partner}}-\\ &\left(\Delta\mathbf{\hat{T}}^{\text{self}}\right)^{-1}\;\mathbf{\hat{T}}_{-1}^{\text{self}\to\text{partner}}\;\Delta\mathbf{\hat{T}}^{\text{partner}}\big)\end{split}$ (16)
Here, $d_{T}$ is the combined distance function (defined in Eq. 8) applied to the predicted pairwise transform $\mathbf{\hat{T}}^{\text{self}\to\text{partner}}$ and the kinematically propagated transform derived from the prediction at the previous token step $\mathbf{\hat{T}}^{\text{self}\to\text{partner}}_{-1}$.
Inference. During inference, DFoT iteratively denoises the sequence $\mathbf{M}^{(\tau)}$ from $\mathbf{\tau}=1\to 0$ using the learned denoiser $\mathbf{f}_{\mathbf{\phi}}$, producing the final predicted clean token sequence $\hat{\mathbf{M}}_{\mathbf{0}}$. The final reconstructed physical motion sequence $\hat{\mathbf{X}}$ is then generated by the VQ-VAE decoder using the predicted latent pose sequence $\mathbf{\hat{Z}}$ and the predicted canonical temporal sequence $\Delta\mathbf{\hat{T}}^{\text{can}}$ as conditioning:
$\hat{\mathbf{X}}\;=\;\mathrm{Dec}(\mathbf{\hat{Z}}\;|\;\beta,\;\Delta\mathbf{\hat{T}}^{\text{can}})$ (17)
We use $\mathbf{\hat{Z}}$ to represent the sequence of all predicted $\hat{\mathbf{z}}$ tokens, and $\Delta\mathbf{\hat{T}}^{\text{can}}$ to represent the full sequence of the intra-agent temporal transforms, extracted from $\hat{\mathbf{M}}_{\mathbf{0}}$.The body shape $\beta$ is constant conditioning over time respectively for each agent.
3.4 Sampling Strategies
Our framework supports multiple sampling modes for multi-agent motion generation at inference time.
Partner In-painting. Given the full sequence of Agent B, the model reconstructs or completes the motion of Agent A:
$P(A_{\text{0:L}}\mid B_{\text{0:L}}).$ (18)
Partner Prediction. We forecast the future of a single agent conditioned on both agents’ past motion:
$P(A_{t:L}\mid A_{0:t-1},B_{0:t-1}).$ (19)
Joint Future Prediction. All agents’ future motion is jointly generated from a single distribution, ensuring coordinated predictions:
$P(A_{t:t+L},B_{t:t+L}\mid A_{0:t-1},B_{0:t-1}).$ (20)
Joint future prediction preserves spatial and temporal correlations and naturally extends to $n\geq 2$ agents.
Agentic Motion Sampling. A scalable approach with per-agent inference, where each agent’s behavior is generated while conditioning on other agents.
Synchronous (Parallel). All agents generate motion at time $t$ in parallel:
$P(A_{t}\mid A_{0:t-1},B_{0:t-1}),\ P(B_{t}\mid A_{0:t-1},B_{0:t-1}).$ (21)
Asynchronous (Turn-Taking). Agents generate motion sequentially, enabling reactive behaviors:
$P(A_{t}\mid A_{0:t-1},B_{0:t-1}),\ P(B_{t}\mid A_{0:t},B_{0:t-1}).$ (22)
Ultra-long Motion Generation. We adopt an autoregressive windowed strategy for continuous sequence generation. The sequence is decomposed into overlapping segments with window size $W$ and overlap $O$, yielding a prediction stride of S=WOS=W-O. At each iteration $k$, the model $f_{\theta}$ predicts a new segment $\tilde{z}_{k\cdot S:k\cdot S+W-1}$ conditioned on the final $O$ frames of the previously generated segment:
$\tilde{z}_{k\cdot S:k\cdot S+W-1}=f_{\theta}\left(z_{(k-1)\cdot S:(k\cdot S)-1}\right)$ (23)
The retained overlap ensures temporal continuity between adjacent windows, allowing the model to generate arbitrarily long and temporally consistent motion sequences.
4 Experiments
Evaluation Metrics. We evaluate our model using standard metrics capturing positional accuracy, motion quality, physical plausibility, and multi-agent coordination: Fréchet Distance (FD) assesses distributional similarity between generated and real motions; Diversity (DIV) measures sample-level per-frame variance to ensure the model avoids mode collapse; Foot Skating (FS) quantifies foot sliding using average skating velocity on ground contact; Interpenetration (IP) uses capsule proxies to detect and measure penetration depth between body parts; Motion Interaction (MI) measures correlation between multiple agents by computing the difference between ground-truth and predicted correlations of their joint positions; Mean Per Joint Position Error (MPJPE) and Mean Per Joint Velocity Error (MPJVE) measure average Euclidean distance between predicted and ground-truth joint positions and velocities, computed as the minimum over 10 generated samples for those methods that can generate multiple samples.
Method In-Painting
FD DIV MI FS IP MPJPE MPJVE
ReMoCap RS 42.151 24.394 0.406 0.303 0.522 1.772 0.041
NN 28.284 - 0.361 0.355 0.610 1.559 0.0381
ReMoS [ghosh2024remos] 0.002 0.000 0.003 0.469 0.162 0.026 0.012
Ours 0.029 0.028 0.000 0.513 0.176 0.074 0.012
DD100 RS 9.73 24.87 0.46 0.49 0.45 1.51 0.03
NN 4.75 - 0.30 0.34 0.69 1.58 0.04
Duolando [siyao2024duolando] 18.18 0.00 0.17 1.88 0.56 1.68 0.07
Ours 0.05 0.13 0.07 0.58 0.12 0.11 0.01
Table 2: Evaluation of in-painting. We evaluate our model on the in-painting task on DD100 and ReMoS datasets against two baselines (RN, NN) and, for each dataset, against an existing model. Dataset name in olive.
Data. We evaluate our method across a diverse set of motion datasets spanning contact sports, dance, and everyday interactions. DuoBoX [cen2025readytoreact] captures high-contact athletic motion; ReMoCap [ghosh2024remos], DD100 [siyao2024duolando], and Embody 3D [mclean2025embody], AMASS, and Inter-X [xu2024inter] represent everyday multi-person interactions. AMASS primarily contains single-person motion, Embody 3D includes one to four interacting people, and the remaining datasets consist of dyadic interactions. In the sup. mat., we provide summary statistics for all datasets.
Implementation Details. For training the 1–4 agent model on Embody 3D containing varying numbers of agents (2–4), we randomly mask $p$ agents during training. This strategy allows the model to generalize across scenes with different agent counts without the need to train separate models for each configuration. We use Viser library to visualize our results [yi2025viser]. Please see sup. mat. for more details.
4.1 Baselines
Classic Baselines: We consider two classic baselines: (1) Nearest Neighbor (NN), which retrieves the most similar sequence from the training set based on the motion token sequence; and (2) Random Sample, which randomly selects a sequence from the training distribution.
SoTA Methods: We compare against the following state-of-the-art methods: (1) Duolando w/o music [siyao2024duolando]: An autoregressive inpainting approach that generates the follower’s motion from the leader’s past motion using look-ahead attention on the leader’s future trajectory. We re-train Duolando without music conditioning. (2) ReMos [ghosh2024remos]: An inpainting method that generates one partner’s motion conditioned on the complete future motion of the other partner. (3) Ready2React (R2R) [cen2025readytoreact]: Handles both partner prediction and dyadic motion generation. For partner prediction, it uses both agents’ past motion to predict one agent’s future. For dyadic prediction, it generates both agents’ motion in an agentic manner from their shared motion history. We use the joint mapping provided in [cen2025readytoreact] to map the MOTIVE skeleton to SMPL-X for evaluation. For a fair comparison with R2R, we also use four frames as past motion when sampling from our model.
For fairness, we train a version of our model on each dataset under similar conditions as the baseline model.
Method Partner Motion Prediction Dyadic Motion Prediction
FD DIV MI FS IP MPJPE MPJVE FD DIV MI FS IP MPJPE MPJVE
DuoBox RS 0.310 5.176 0.156 0.277 0.350 0.119 5.413 0.119 0.342 0.394 0.735 0.028
@@ 4.1 Baselines
Ours 0.310 1.764 0.012 0.383 0.399 0.599 0.029 0.118 5.622 0.000 0.407 0.101 0.714 0.034
Table 3: Evaluation on DuoBox. Our model is evaluated on Partner Motion Prediction and Dyadic Motion Prediction, benchmarked against two classical baselines (RS, NN) and the SOTA method (R2R). We include an ablation on Dyadic Prediction to analyze the effect of offsetting denoising steps for synthesizing Turn-Taking ($T$) variations. Dataset name in olive.
Dataset Joint Future Prediction
FID DIV MI FS IP MPJPE MPJVE
InterX 0.210 2.911 0.130 0.074 0.093 0.475 0.013
Embody3D 2 people 1.409 7.573 0.072 0.312 0.023 0.712 0.011
3 people 0.477 8.039 0.008 0.283 0.046 0.612 0.010
4 people 1.032 5.825 0.160 0.292 0.044 0.744 0.011
Table 4: Evaluation on Social Interaction Dataset. Our model is evaluated on joint future prediction on InterX and Embody3D. For Embody3D, we train a single model on the full multi-person set (2–4 people) and report metrics separately on the 2-, 3-, and 4-person subsets.
@@ 4.2 Quantitative Results
On the in-painting task in Table 2, our method performs comparably to existing approaches, with varying strengths across datasets. On ReMoCap, our results are largely on-par with the dedicated ReMoS model [ghosh2024remos]: FS, IP, and MPJPE remain close (FS 0.513 vs. 0.469, IP 0.176 vs. 0.162, MPJPE 0.074 vs. 0.026). ReMoS achieves exceptionally low FD; however, this comes at the cost of notably low diversity (DIV 0.028 vs. 0.000). Our model, in contrast, achieves slightly higher FD but produces more diverse outputs, offering a more balanced trade-off between realism and variation. The NN and RS baselines achieve good results in foot sliding which is to be expected while the interaction metrics (MI and IP) and motion realism of two people does not compete with our model. On the DD100 test set, we obtain the lowest (best) FD score (0.05 vs. 18.18 for Duolando (trained without music) [siyao2024duolando]) and competitive performance across other metrics. The relatively high FD score of DD100 suggests that the music signal is an important resource for Duolando. In both cases, our model outperforms previous works in terms of motion interaction indicating it’s ability to model realistic human interaction.
@@
In Table 3 we compare our model to Ready-to-React [cen2025readytoreact] on the DuoBox test set for both partner-motion prediction and dyadic-motion generation. On the partner-prediction task, our method achieves competitive results with R2R. For instance, the Interpenetration errors (IP: 0.399 vs. 0.309) are close, and we match R2R exactly on MPJVE (0.029 for both). While R2R shows lower FD and DIV in the prediction setting, our model attains superior FD and DIV scores in the more challenging dyadic motion prediction task, indicating better global motion stability and diversity when generating full two-person interactions. Moreover, our approach produces the lowest Interpenetration and Foot Skating errors in the dyadic setting, underscoring the physical plausibility and interaction fidelity of our generated motions.
@@
Overall, the results underscore the key advantage of our approach: despite not being tailored to any single task or dataset, our method remains competitive on all metrics and particularly strong on those reflecting realism and interaction quality. This balance of generality and performance demonstrates the value of our unified formulation.
@@
Ablations. To assess our architectural design choices, we ablate key components and analyze their impact on generation quality in Table 5. No VQ-VAE (Raw Joints and $\mathbf{T}^{\text{can}\to\text{root}}$ ) assesses the contribution of motion quantization, we replace the discrete VQ-VAE embeddings with continuous body parameters and canonical-to-root transformations. Table 5 shows that without VQ-VAE, the overall model performance declines significantly. Importance of $\mathbf{T}^{\text{self}\to\text{partner}}_{t}$: We remove the relative transformation $\mathbf{T}^{\text{self}\to\text{partner}}_{t}$ from the motion tokens to evaluate the impact of explicit inter-agent spatial alignment. Its removal eliminates inter-agent spatial encoding, leading to significantly worse model performance. As seen in Table 5, MPJPE rose by 3.6% (from 0.62 to 0.65), and penetration artifacts worsened tenfold (0.002 to 0.025). This demonstrates that explicit spatial alignment modeling is essential for plausible and coordinated interactions.
Method Joint Future Prediction
FID DIV MI FS IP MPJPE MPJVE
DuoBox w/o $\mathbf{T}^{\text{self}\to\text{partner}}$ 4.476 5.174 0.159 0.370 0.354 0.678 0.006
w/o VQVAE 0.190 11.367 0.255 5.688 0.057 1.128 0.011
MAGNet (Ours) 0.052 9.572 0.124 0.423 0.116 0.641 0.006
Table 5: Ablation of model components on DuoBox to evaluate the impact of motion quantization and inter-agent spatial alignment on generation quality. Dataset name in olive
@@
We further analyze in  Table 3, the effect of varying the offset interval in our agentic turn-taking sampling schedule, which determines when subsequent tokens begin denoising. This ablation studies how different offset steps influence the generation. We observe that smaller offset steps lead to more dynamic and responsive interactions, whereas larger offsets promote smoother but less reactive motion, as the agent waits until the other agent is completely denoised. Please see the supplemental video for results on more inference strategies.
Method Dyadic Motion Generation
FD DIV MI FS IP MPJPE MPJVE
Embody3D 4 people w/o HG 1.032 5.825 0.160 0.292 0.044 0.744 0.011
w/ HG 0.934 5.192 0.154 0.304 0.045 0.732 0.011
w/ SHG 1.123 2.910 0.205 0.237 0.030 0.736 0.011
w/ PHG 0.964 4.359 0.179 0.265 0.034 0.721 0.010
Table 6: Effect of History Guidance on Embody3D (4 people). We report joint future prediction metrics with and without history guidance. Dataset name is shown in olive.
Effect of History Guidance.
We investigate whether guidance improves joint future prediction on Embody3D (4-people). Inspired by History Guidance (HG) [song2025historyguidedvideodiffusion], we also explore two variants that condition guidance on different temporal contexts: Self History Guidance (SHG), which uses an agent’s own past, and Partner History Guidance (PHG), which uses the interaction partner’s past (see supplemental for details).
As shown in Table 5, all guidance variants yield modest improvements over the no-guidance baseline on selected metrics, although the overall performance differences remain small, indicating that history guidance offers limited but consistent benefits in this setting. Notably, different guidance strategies emphasize distinct aspects of motion quality: standard HG primarily improves global distributional alignment (FD, MI), SHG favors self-consistency and interaction smoothness (FS, IP), and PHG enhances partner-aware kinematic accuracy (MPJPE, MPJVE). These results suggest that history guidance introduces targeted inductive biases rather than delivering a uniform performance gain across metrics.

### v2

@@ 4.1 Baselines
RS 0.310 5.176 0.156 0.277 0.350 0.119 5.413 0.119 0.342 0.394 0.735 0.028
@@
Ours 0.067 0.669 0.020 0.270 0.019 0.218 0.016 0.118 5.622 0.000 0.407 0.101 0.714 0.034
Table 3: Evaluation on DuoBox. Our model is evaluated on Partner Motion Prediction and Dyadic Motion Prediction, benchmarked against two classical baselines (RS, NN) and the SOTA method (R2R). We include an ablation on Dyadic Prediction to analyze the effect of offsetting denoising steps for synthesizing Turn-Taking ($T$) variations.
@@ 4.2 Quantitative Results
On the in-painting task in Table 2, our method performs comparably to existing approaches, with varying strengths across datasets. On ReMoCap, our results are largely on-par with the dedicated ReMoS model [9]: FS, IP, and MPJPE remain close (FS 0.513 vs. 0.469, IP 0.176 vs. 0.162, MPJPE 0.074 vs. 0.026). ReMoS achieves exceptionally low FD; however, this comes at the cost of notably low diversity (DIV 0.028 vs. 0.000). Our model, in contrast, achieves slightly higher FD but produces more diverse outputs, offering a more balanced trade-off between realism and variation. The NN and RS baselines achieve good results in foot sliding which is to be expected while the interaction metrics (MI and IP) and motion realism of two people does not compete with our model. On the DD100 test set, we obtain the lowest (best) FD score (0.05 vs. 18.18 for Duolando (trained without music) [26]) and competitive performance across other metrics. The relatively high FD score of DD100 suggests that the music signal is an important resource for Duolando. In both cases, our model outperforms previous works in terms of motion interaction indicating it’s ability to model realistic human interaction.
Partner Prediction.
In Table 3 we compare our model to Ready-to-React [2] on the DuoBox test set. On the partner-prediction task, our method outperforms R2R on nearly all metrics, achieving the best FD (0.067 vs. 0.181), MI (0.020 vs. 0.071), IP (0.019 vs. 0.309), MPJPE (0.218 vs. 0.580), and MPJVE (0.016 vs. 0.029). These gains suggest that diffusion forcing over multi-agent tokens is effective even in the simpler single-agent prediction setting, where it’s only conditioned on past motion.
Dyadic Prediction.
@@
On the more challenging dyadic motion generation task, where both agents are generated, our model attains superior FD and DIV scores, indicating better global motion stability and diversity when generating full two-person interactions. Moreover, our approach produces the lowest Interpenetration and Foot Skating errors in the dyadic setting, underscoring the physical plausibility and interaction fidelity of our generated motions.
Polyadic Future Prediction.
A key advantage of our architecture is that a single trained model naturally extends to any number of agents without modification. To our knowledge, no existing learned method supports generation beyond two agents; we therefore evaluate against the only universally applicable references: NN and RS. Table 4 reports results on Embody3D as N increases from 2 to 4 using the same model. Our method achieves the best FD across all group sizes and maintains the lowest MPJPE and foot skating (FS), with the advantage widening at N=4—where NN and RS see sharp FD degradation due to data sparsity, our model degrades gracefully ($0.298\rightarrow 0.590$ vs. $0.659\rightarrow 2.143$for NN). NN achieves low MI and MPJVE by construction, as it retrieves real sequences, but this does not extend to realistic generation for larger groups. RS yields high diversity but lacks inter-agent coherence. Our method’s lower diversity relative to RS reflects a natural consequence of generating coordinated group motion rather than independent samples, as confirmed qualitatively in Figure 6.
[…]
2 RS 0.885 7.408 0.052 0.332 0.007 0.668 0.008
NN 0.659 0.065 0.471 0.224 0.646 0.007
Ours 0.298 6.219 0.093 0.207 0.044 0.561 0.011
3 RS 0.431 7.948 0.011 0.284 0.036 0.617 0.006
NN 0.566 0.019 0.298 0.700 0.539 0.006
Ours 0.293 6.313 0.004 0.257 0.043 0.526 0.011
4 RS 2.039 12.945 0.007 0.310 0.054 0.894 0.009
NN 2.143 0.001 0.363 0.573 0.880 0.009
Ours 0.590 3.018 0.058 0.237 0.050 0.546 0.009
Table 4: Polyadic quantitative evaluation on Embody3D. Best results are bolded per NN and per metric.
Figure 7: Cumulative FD over generation length.
Ultra-Long Generation.
Our method significantly outperforms R2R over 1800 frames on Duobox with 4 frames as context, achieving an FD of 2.781 vs. 24.810, MI of 0.021 vs. 0.062, and IP of 0.116 vs. 0.345. Figure 7 shows that R2R’s FD accumulates exponentially, while MAGNet maintains superior stability, demonstrating robustness to temporal drift.
@@
Overall, these results underscore the key advantage of our approach: despite not being tailored to any single task or dataset, our method remains competitive across all metrics and particularly strong on those reflecting long-horizon coherence, high realism, and interaction quality. This balance of generality and performance demonstrates the value of our unified formulation.
Ablations.
@@
To assess our architectural design choices, we ablate key components and analyze their impact on generation quality in Table 5. No VQ-VAE (Raw Joints and $\mathbf{T}^{\text{can}\to\text{root}}$ ) assesses the contribution of motion quantization, we replace the discrete VQ-VAE embeddings with continuous body parameters and canonical-to-root transformations. Table 5 shows that without VQ-VAE, the overall model performance declines significantly. Importance of $\mathbf{T}^{\text{self}\to\text{partner}}_{t}$: To evaluate the impact of explicit inter-agent spatial alignment, we remove the relative transformation $\mathbf{T}^{\text{self}\to\text{partner}}_{t}$ from the motion tokens. The removal of this inter-agent spatial encoding leads to a significant degradation in model performance. As shown in Table 5, FD and MPJPE deteriorate (rising from 0.052 to 0.126 and 0.641 to 0.665, respectively), while penetration artifacts nearly triple, increasing from 0.116 to 0.342. These results demonstrate that explicit spatial alignment modeling is essential for synthesizing physically plausible and coordinated interactions.
[…]
w/o VQ-VAE 4.476 5.174 0.255 5.682 0.057 1.128 0.200
w/o $\mathbf{T}^{\text{self}\to\text{partner}}$ 0.126 10.761 0.128 0.367 0.342 0.665 0.032
MAGNet (Ours) 0.052 9.572 0.124 0.423 0.116 0.641 0.034
Table 5: the VQ-VAE causes catastrophic foot skating (FS: 5.682) and FD degradation ($86\times$), indicating that discrete tokenization regularizes motion quality. Removing $\mathbf{T}^{\text{self}\to\text{partner}}$ nearly triples interpenetration (IP: 0.116 $\to$ 0.342), confirming that explicit inter-agent spatial encoding is critical for coordination.
@@
We further analyze in  Table 3, the effect of varying the offset interval in our agentic turn-taking sampling schedule, which determines when subsequent tokens begin denoising. This ablation studies how different offset steps influence the generation. We observe that smaller offset steps lead to more dynamic and responsive interactions, whereas larger offsets promote smoother but less reactive motion, as the agent waits until the other agent is completely denoised. Please see the supplementary video for results on more inference strategies.